If the sun expanded to a radius 200 times it's present radius, what would be it's average density be in g/cm^3
1
Expert's answer
2020-06-17T09:35:39-0400

Radius of expanded sun is "R_e=200R"

Mass of Sun is "M=1.989 \u00d7 10^{33}" gram.

Volume of expanded Sun will be

"V_e=\frac{4}{3}\pi R^3_e"

Density will be,

"\rho_e=\frac{M}{V_e}=\frac{3M}{4\pi R^3_e}\\
\rho_e=\frac{1}{200^3}\rho"

where, "\rho" is the density of Sun before expansion begin. Thus,

"\rho_e=1.76\times 10^{-7}\:g/cm^3"